### v2

@@ Acknowledgments
(S(25)) $\left|\left(\nabla_{\Theta}^{\mathsf{q}}E\right)_{l}-\left.\frac{\delta\mathcal{E}}{\delta{\theta}}\right|_{\theta(t)}(lh)\right|$ $=|\hat{\mathsf{q}}_{l}^{\top}\partial_{\theta}f(z(lh),\mathsf{\theta}_{l})-p^{\top}(lh)\partial_{\theta}f(z(lh),\mathsf{\theta}_{l})|$
@@
where we used the boundedness of $z$, the regularity of the first order derivative of $f$ and Lemma S2.6.
@@
Lemma S2.13 (Control of Term II in (25)).
@@
(S(26)) $|L\mathsf{T}\nabla_{\Theta}^{\mathsf{p}}E-L\nabla_{\Theta}^{\mathsf{q}}E|=\mathcal{O}(h^{2})\ ,$
@@
Appendix S3 Proof of Proposition 3.3 and Theorem 3.4
This section is dedicated to 2-stage ERK ODE-nets. We sketch the proof of Proposition 3.3 and Theorem 3.4 in Section S3.1. The proof for Theorem 3.4 calls for some a-priori estimates and these technical controls are provided in Section S3.2.
S3.1 Main proof of Proposition 3.3 and Theorem 3.4
Proof S3.1 (Proof of Proposition 3.3).
We first note that auto-differentiation is obtained using chain rule [17, 27], and thus:
(S(27)) $\partial_{\mathsf{z}_{l}}E=(\partial_{\mathsf{z}_{l+1}}E)(\partial_{\mathsf{z}_{l}}\mathsf{z}_{l+1})\,.$
@@
$\partial_{\mathsf{z}_{L}}E=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})=\mathsf{p}_{L}^{\top}\ .$
@@
(S(28)) $\partial_{\mathsf{z}_{l}}\mathsf{z}_{l+\alpha}=I+\alpha h\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,.$
@@
where in the first equality we applied chain rule, and in the second equality we plugged in (S(28)). Plugging this in (S(27)):
@@
Proof S3.2 (Proof of Theorem 3.4).
@@
where in both we used triangular inequality. If we can provide upper bound of $p$, $z$, $\mathsf{z}$ and control the difference terms using $\mathcal{O}(h^{2})$, the proof is complete. The boundedness of $p$, $z$, $\mathsf{z}$ are collected in Lemma S3.3, and the difference terms are controlled in Corollary S3.6 and S3.13 respectively.
S3.2 A-priori estimates for 2-stage ERK
We record two very useful a-priori results for 2-stage ERK ODE-nets here. The first result on boundedness resembles Lemma S2.1. It is once again an analog to  [9] and its proof is a straightforward extension that has been omitted here.
Lemma S3.3.
@@
(S(30)) $|z((l+\alpha)h)-\mathsf{z}_{l+\alpha}|=\mathcal{O}(h^{2})\,.$
@@
(S(31)) $|\partial_{\theta}f(\mathsf{z}_{l},\theta_{l})-\partial_{\theta}f(z(lh),\theta_{l})|=\mathcal{O}(h^{2})\,$,
@@
(S(32)) $|\partial_{\theta}f(\mathsf{z}_{l+\alpha},\theta_{l+\alpha})-\partial_{\theta}f(z((l+\alpha)h),\theta_{l+\alpha})|=\mathcal{O}(h^{2})\,.$
Proof S3.7.
Both estimate follows from Lemma S3.4 and third order continuous differentiability and bounded derivatives assumptions on $f$.
@@
Proof S3.8 (Proof of Proposition 2.4).
@@
where we applied triangular inequality. The two terms are controlled in Lemma S3.9 and Lemma S3.11.
@@
(S(33)) $|\hat{\mathsf{p}}_{l}-{\mathsf{p}}_{l}|=\mathcal{O}(h^{2})\,$,
@@
$=|(g(z(1))-y(x))\nabla g(z(1))-(g({\mathsf{z}}_{L})-y(x))\nabla g({\mathsf{z}}_{L})|$
@@
where we used Lemma S3.4, the Lipschitz continuity assumption on $g$ and $\nabla g$, and the boundedness of $\mathsf{z}_{L}$. For $l=0,1,\dots,L-1$ we call (18) to get
@@
(S(35)) $\frac{d^{2}p^{\top}}{dt^{2}}$ $=-\frac{d}{dt}(p^{\top}\partial_{z}f(z(t),\theta(t)))$
@@
(S(36)) $\ \ \ \ |p^{\top}((l+1)h)+\frac{h}{2}p^{\top}((l+1)h)\partial_{z}f(z((l+1)h),\mathsf{\theta}_{l+1})$
@@
where we used a linear approximation to $\frac{d}{dt}(\partial_{z}f(z(t),\theta(t)))$ in the second equality, triangular inequality in the first inequality, and second order Taylor approximation, (12) and (S(35)) in the last line.
@@
where in the first inequality we used triangular inequality and (S(36)), in the second equality we subtracted and then added $\hat{\mathsf{p}}_{l+1}^{\top}$ in order to create the difference $\Delta_{l+1}^{\hat{\mathsf{p}}}$, in the third inequality we again applied the triangular inequality, and in the last inequality we utilized Lemma S3.4.
To bound the last term of (S(37)), we compute:
@@
It is easy to see that the two factors in Term I are both $\frac{d}{dt}\partial_{z}f(z(lh),\theta(lh))+\mathcal{O}(h)$, so they cancel out, leaving another $\mathcal{O}(h)$, making $\text{Term I}=\mathcal{O}(h^{3})$. To analyze Term II, we also notice that the difference term contribute one extra $\mathcal{O}(h)$, making the whole term $\text{Term II}=\mathcal{O}(h^{3})$. Details are heavily repeated and thus not shown, and Lemma S3.3 is called for the boundedness of $\hat{\mathsf{p}}_{l+1}$ and $z$.
Now plugging these controls back into (S(37)), we get
@@
(S(40)) $|\mathsf{q}_{l+\alpha}-p((l+\alpha)h)|$
@@
Appendix S4 Additional Figures
@@
Figure S1: Gradient computation for the Midpoint Method. The left panel compares $L\nabla_{\Theta}E$ (green) with $\mathsf{NN}_{\Theta}$ formulated in (6) with $\frac{\delta\mathcal{E}}{\delta\theta}$ (blue). $L\nabla_{\Theta}E$ oscillates wildly, recovering the true gradient only at half-steps. Our modification (red) yields gradients that match the ground-truth. The panel on the right shows convergence with respect to $h$. The $L_{\infty}$ norm of the gradient difference stays as a constant even when $h$ is small, while our modified gradient converges to the ground-truth with a rate of $h^{1.70}$, confirming the theorem. The gradient from using forward Euler obtains the rate of $0.98$ (red and blue).
@@
Figure S2: Gradient computation for the Ralston Method. The left panel compares $L\nabla_{\Theta}E$ (green) with $\mathsf{NN}_{\Theta}$ being defined as the Ralston method with $\frac{\delta\mathcal{E}}{\delta\theta}$ (blue). The output of auto-differentiation $L\nabla_{\Theta}E$ oscillates wildly, and is far away from the true gradient. Our modification (red) yields gradients that match the ground-truth. The right panel shows convergence. The $L_{\infty}$ error of the output of auto differentiation stays as a constant. The convergence rates for the modified gradient and that computed from forward Euler type discretization are $1.84$ and $0.98$ respectively (red and blue).
@@
Figure S3: For Midpoint Method. The top row shows the training loss and test error in epochs. The middle row compares $\Theta$ outputs from vanilla auto-differentiation method and our modified gradients. The bottom plot presents the final actual reconstruction.
@@
Figure S4: For Ralston Method. The top row shows the training loss and test error in epochs. The middle row compares $\Theta$ outputs from vanilla auto-differentiation method and our modified gradients. The bottom plot presents the final actual reconstruction.
@@
Figure S5: For Midpoint Method. The figure compares reconstructions of a spiral learned using auto-differentiation gradients (red) and our modified gradients (blue). Our modification achieves results that align more closely with the ground truth (green).
@@
Figure S6: For Ralston Method. The figure compares reconstructions of a spiral learned using auto-differentiation gradients (red) and our modified gradients (blue). Our modification achieves results that align more closely with the ground truth (green).

### v3

@@ Acknowledgments
(SM(25)) $\left|\left(\nabla_{\Theta}^{\mathsf{q}}E\right)_{l}-\left.\frac{\delta\mathcal{E}}{\delta{\theta}}\right|_{\theta(t)}(lh)\right|$ $=|\hat{\mathsf{q}}_{l}^{\top}\partial_{\theta}f(z(lh),\mathsf{\theta}_{l})-p^{\top}(lh)\partial_{\theta}f(z(lh),\mathsf{\theta}_{l})|$
@@
where we used the boundedness of $z$, the regularity of the first order derivative of $f$ and Lemma SM2.6.
@@
Lemma SM2.13 (Control of Term II in (25)).
@@
(SM(26)) $|L\mathsf{T}\nabla_{\Theta}^{\mathsf{p}}E-L\nabla_{\Theta}^{\mathsf{q}}E|=\mathcal{O}(h^{2})\ ,$
@@
Appendix SM3 Proof of Proposition 3.3 and Theorem 3.4
This section is dedicated to 2-stage ERK ODE-nets. We sketch the proof of Proposition 3.3 and Theorem 3.4 in Section SM3.1. The proof for Theorem 3.4 calls for some a-priori estimates and these technical controls are provided in Section SM3.2.
SM3.1 Main proof of Proposition 3.3 and Theorem 3.4
Proof SM3.1 (Proof of Proposition 3.3).
We first note that auto-differentiation is obtained using chain rule [Linnainmaa1976, Rumelhart1986], and thus:
(SM(27)) $\partial_{\mathsf{z}_{l}}E=(\partial_{\mathsf{z}_{l+1}}E)(\partial_{\mathsf{z}_{l}}\mathsf{z}_{l+1})\,.$
@@
$\partial_{\mathsf{z}_{L}}E=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}=\mathsf{p}_{L}^{\top}\ .$
@@
(SM(28)) $\partial_{\mathsf{z}_{l}}\mathsf{z}_{l+\alpha}=I+\alpha h\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,.$
@@
where in the first equality we applied chain rule, and in the second equality we plugged in (SM(28)). Plugging this in (SM(27)):
@@
Proof SM3.2 (Proof of Theorem 3.4).
@@
where in both we used triangular inequality. If we can provide upper bound of $p$, $z$, $\mathsf{z}$ and control the difference terms using $\mathcal{O}(h^{2})$, the proof is complete. The boundedness of $p$, $z$, $\mathsf{z}$ are collected in Lemma SM3.3, and the difference terms are controlled in Corollary SM3.6 and SM3.13 respectively.
SM3.2 A-priori estimates for 2-stage ERK
We record two very useful a-priori results for 2-stage ERK ODE-nets here. The first result on boundedness resembles Lemma SM2.1. It is once again an analog to  [DiChLiWr:2022overparameterization, Lemmas 14 and 24] and its proof is a straightforward extension that has been omitted here.
Lemma SM3.3.
@@
(SM(30)) $|z((l+\alpha)h)-\mathsf{z}_{l+\alpha}|=\mathcal{O}(h^{2})\,.$
@@
(SM(31)) $|\partial_{\theta}f(\mathsf{z}_{l},\theta_{l})-\partial_{\theta}f(z(lh),\theta_{l})|=\mathcal{O}(h^{2})\,$,
@@
(SM(32)) $|\partial_{\theta}f(\mathsf{z}_{l+\alpha},\theta_{l+\alpha})-\partial_{\theta}f(z((l+\alpha)h),\theta_{l+\alpha})|=\mathcal{O}(h^{2})\,.$
Proof SM3.7.
Both estimate follows from Lemma SM3.4 and third order continuous differentiability and bounded derivatives assumptions on $f$.
@@
Proof SM3.8 (Proof of Proposition 2.4).
@@
where we applied triangular inequality. The two terms are controlled in Lemma SM3.9 and Lemma SM3.11.
@@
(SM(33)) $|\hat{\mathsf{p}}_{l}-{\mathsf{p}}_{l}|=\mathcal{O}(h^{2})\,$,
@@
$=|\langle(g(z(1))-y(x)),\nabla g(z(1))\rangle_{\mu}-\langle(g({\mathsf{z}}_{L})-y(x)),\nabla g({\mathsf{z}}_{L})\rangle_{\mu}|$
@@
where we used Lemma SM3.4, the Lipschitz continuity assumption on $g$ and $\nabla g$, and the boundedness of $\mathsf{z}_{L}$. For $l=0,1,\dots,L-1$ we call (18) to get
@@
(SM(35)) $\frac{d^{2}p^{\top}}{dt^{2}}$ $=-\frac{d}{dt}(p^{\top}\partial_{z}f(z(t),\theta(t)))$
@@
(SM(36)) $\ \ \ \ |p^{\top}((l+1)h)+\frac{h}{2}p^{\top}((l+1)h)\partial_{z}f(z((l+1)h),\mathsf{\theta}_{l+1})$
@@
where we used a linear approximation to $\frac{d}{dt}(\partial_{z}f(z(t),\theta(t)))$ in the second equality, triangular inequality in the first inequality, and second order Taylor approximation, (12) and (SM(35)) in the last line.
@@
where in the first inequality we used triangular inequality and (SM(36)), in the second equality we subtracted and then added $\hat{\mathsf{p}}_{l+1}^{\top}$ in order to create the difference $\Delta_{l+1}^{\hat{\mathsf{p}}}$, in the third inequality we again applied the triangular inequality, and in the last inequality we utilized Lemma SM3.4.
To bound the last term of (SM(37)), we compute:
@@
It is easy to see that the two factors in Term I are both $\frac{d}{dt}\partial_{z}f(z(lh),\theta(lh))+\mathcal{O}(h)$, so they cancel out, leaving another $\mathcal{O}(h)$, making $\text{Term I}=\mathcal{O}(h^{3})$. To analyze Term II, we also notice that the difference term contribute one extra $\mathcal{O}(h)$, making the whole term $\text{Term II}=\mathcal{O}(h^{3})$. Details are heavily repeated and thus not shown, and Lemma SM3.3 is called for the boundedness of $\hat{\mathsf{p}}_{l+1}$ and $z$.
Now plugging these controls back into (SM(37)), we get
@@
(SM(40)) $|\mathsf{q}_{l+\alpha}-p((l+\alpha)h)|$
@@
Appendix SM4 Additional Figures
@@
Figure SM1: Gradient computation for the Midpoint Method. The left panel compares $L\nabla_{\Theta}E$ (green) with $\mathsf{NN}_{\Theta}$ formulated in (6) with $\frac{\delta\mathcal{E}}{\delta\theta}$ (blue). $L\nabla_{\Theta}E$ oscillates wildly, recovering the true gradient only at half-steps. Our modification (red) yields gradients that match the ground-truth. The panel on the right shows convergence with respect to $h$. The $L_{\infty}$ norm of the gradient difference stays as a constant even when $h$ is small, while our modified gradient converges to the ground-truth with a rate of $h^{1.70}$, confirming the theorem. The gradient from using forward Euler obtains the rate of $0.98$ (red and blue).
@@
Figure SM2: Gradient computation for the Ralston Method. The left panel compares $L\nabla_{\Theta}E$ (green) with $\mathsf{NN}_{\Theta}$ being defined as the Ralston method with $\frac{\delta\mathcal{E}}{\delta\theta}$ (blue). The output of auto-differentiation $L\nabla_{\Theta}E$ oscillates wildly, and is far away from the true gradient. Our modification (red) yields gradients that match the ground-truth. The right panel shows convergence. The $L_{\infty}$ error of the output of auto differentiation stays as a constant. The convergence rates for the modified gradient and that computed from forward Euler type discretization are $1.84$ and $0.98$ respectively (red and blue).
@@
Figure SM3: For Midpoint method. The top row shows the training loss and test error over epochs. The middle row compares $\Theta$ outputs from vanilla (unmodified) auto-differentiation and our modified gradients, and the bottom plot presents the final reconstruction. Quantitatively, the corrected gradients achieve much lower final errors: training loss decreases from $4.2\times 10^{-4}$ to $2.4\times 10^{-4}$, and test error from $9.0\times 10^{-4}$ to $5.5\times 10^{-4}$. These improvements confirm the effectiveness of the proposed correction.
@@
Figure SM4: For Ralston Method. The top row shows the training loss and test error over epochs. The middle row compares $\Theta$ outputs from vanilla (unmodified) auto-differentiation and our modified gradients, and the bottom plot presents the final reconstruction. Quantitatively, the corrected gradients achieve much lower final errors: training loss decreases from $2.8\times 10^{-3}$ to $2.4\times 10^{-4}$, and test error from $3.0\times 10^{-3}$ to $5.4\times 10^{-4}$. These improvements confirm the effectiveness of the proposed correction.
@@
Figure SM5: For Midpoint Method. The figure compares reconstructions of a spiral learned using auto-differentiation gradients (red) and our modified gradients (blue). Our modification achieves results that align more closely with the ground truth (green). Quantitatively, the trajectory Root Mean Square Error (RMSE) drops from $0.23$ to $0.21$, the final-point error from $0.32$ to $0.25$, and the maximum deviation from $0.32$ to $0.28$. These results confirm that the corrected gradients yield trajectories that closely follow the ground-truth dynamics.
@@
Figure SM6: For Ralston Method. The figure compares reconstructions of a spiral learned using auto-differentiation gradients (red) and our modified gradients (blue). Our modification achieves results that align more closely with the ground truth (green). Quantitatively, the trajectory Root Mean Square Error (RMSE) drops from $0.40$ to $0.23$, the final-point error from $0.79$ to $0.31$, and the maximum deviation from $0.79$ to $0.31$. These results confirm that the corrected gradients yield trajectories that closely follow the ground-truth dynamics.